OpenStudy (anonymous):

Let u = <-6, 1>, v = <-5, 2>. Find -4u + 2v. @ganeshie8 @hero @dan815 @pooja195 @triciaal @loser66 @wio @luigi0210

OpenStudy (zzr0ck3r):

hint: \(-4u=\langle -4(-6),-4(1)\rangle = \langle 24, -4\rangle\)

OpenStudy (anonymous):

i really have no idea wow i feel dumb

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

please help @zzr0ck3r

OpenStudy (usukidoll):

adding vectors is in the form \[u_1+v_1,u_2+v_2,...u_n+v_n \] but before we can do that we need to distribute the -4 on the u vector and distribute the 2 on the v vector.

OpenStudy (anonymous):

<34, -8> <14, 0> <14, 3> <44, -12> @UsukiDoll @ganeshie8 these are my options

OpenStudy (usukidoll):

zzrocker already did the distribute -4 all over the u vector now we have to distribute 2 all over the v vector 2<-5,2>

OpenStudy (usukidoll):

now what's -5 x 2 and 2 x 2

OpenStudy (anonymous):

-10 and 4 @UsukiDoll

OpenStudy (usukidoll):

ok cool... so our 2v is <-10,4> so let's add them together zzrocker already did -4u which was <24,-4> so now we have <24,-4> +<-10,4> \[<u_1,u_2> + <v_1,v_2>\] so we have \[u_1 = 24, u_2 = -4, v_1 = -10, v_2 = 4\] but our final answer has to be in the form \[<u_1+v_1,u_2+v_2> \]

OpenStudy (anonymous):

one of my option is 14,0 is that the answer?

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

u vector \[<u_1,u_2...u_n>\] v vector \[<v_1,v_2,...v_n>\] u+v vector \[u_1+v_1,u_2+v_2+...u_n+v_n\]

OpenStudy (usukidoll):

\[<24-10,-4+4 >\]

OpenStudy (anonymous):

wait so was i right? @UsukiDoll

OpenStudy (usukidoll):

yes .. 24-10 = 14 and -4+4 = 0 your u+v vector is <14,0>

OpenStudy (anonymous):

can u help me with another one? @UsukiDoll

OpenStudy (usukidoll):

sure

OpenStudy (anonymous):

Let u = <-9, 4>, v = <8, -5>. Find u - v. <-1, -1> <-13, 13> <-17, 9> <-4, -4>

OpenStudy (usukidoll):

ok this is similar to the addition vector only we are dealing with subtraction u vector \[<u_1,u_2,...u_n> \] v vector \[<v_1,v_2,...v_n >\] u-v vector \[<u_1-v_1,u_2-v_2,....u_n-v_n>\]

OpenStudy (usukidoll):

so... we just have to match labels since we're not multiplying this time.

OpenStudy (anonymous):

so what do i do next? can u take me step by step with numbers? @UsukiDoll

OpenStudy (usukidoll):

sure

OpenStudy (usukidoll):

let's start with the u vector u=<-9,4> recall that our u vector is in the form of \[u_1,u_2,...u_n \] since there are only 2 terms in our u vector, we have something like \[u=<u_1,u_2>\] Therefore, \[u_1=-9,u_2 = 4\]

OpenStudy (usukidoll):

similarly for the v vector v= <8,-5> there are only two terms in our v vector, so we have something like \[v=<v_1,v_2> \] Therefore, \[v_1 = 8,v_2=-5\]

OpenStudy (anonymous):

so its -1 and -1 rights?

OpenStudy (usukidoll):

now we have to do subtraction, which is finding the u-v vector which is in the form \[u-v=<u_1-v_1,u_2-v_2...u_n-v_n>\]

OpenStudy (usukidoll):

only one of them will be -1. Recall these values \[u_1=-9,u_2 = 4 \] \[v_1 = 8,v_2=-5 \] now our u-v vector is in the form \[u-v=<u_1-v_1,u_2-v_2...u_n-v_n>\] or in this situation just \[u-v=<u_1-v_1,u_2-v_2>\]

OpenStudy (anonymous):

can you help me with another ? @UsukiDoll

OpenStudy (usukidoll):

but did you get the final answer first?

OpenStudy (anonymous):

yes it was -1,-1 @UsukiDoll

OpenStudy (anonymous):

Evaluate the expression. r = <9, -7, -1>, v = <2, 2, -2>, w = <-5, -2, 6> v ⋅ w <-18, 14, -2> -26 1 <-10, -4, -12>

OpenStudy (usukidoll):

the previous answer isn't right \[u-v = <-9-8, 4-(-5)>\] please try again before we can move further.

OpenStudy (usukidoll):

what's -9-8? and what's 4-(-5) (distribute the negative)

OpenStudy (anonymous):

-17,9

OpenStudy (usukidoll):

there we go :)

OpenStudy (usukidoll):

so for the next question we are dealing with dot product

OpenStudy (usukidoll):

\[u \cdot v =<u_1v_1,u_2v_2,...u_nv_n> \]

OpenStudy (anonymous):

i think the answer is <-10, -4, -12> @UsukiDoll

OpenStudy (usukidoll):

only it's just \[v \cdot w \] where v = <2,2,-2 > and w = <-5,-2,6> yeah you're right \[v_1=2,v_2=2,v_3=-2...w_1=-5,w_2=-2,w_3=6\] \[v_1w_1=-10,v_2w_2=-4,v_3w_3=-12\] <-10,-4,-12>

OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. B = 46°, a = 12, b = 11 A = 38.3°, C = 95.7°, c = 8; A = 141.7°, C = 84.3°, c = 8 A = 51.7°, C = 82.3°, c = 8; A = 128.3°, C = 5.7°, c = 8 A = 38.3°, C = 95.7°, c = 15.2; A = 141.7°, C = 84.3°, c = 15.2 A = 51.7°, C = 82.3°, c = 15.2; A = 128.3°, C = 5.7°, c = 1.5 @UsukiDoll

OpenStudy (usukidoll):

I'm with another question atm.